OpenStudy (hari5719):

calculate the mass of 9 moles of CO2 {at mass of He=4}

OpenStudy (hari5719):

@aprehan , @abb0t , @ganeshie8

OpenStudy (hari5719):

ill fan u and give a medal pls answer

OpenStudy (hari5719):

@aprehan , @anonymous_user

OpenStudy (hari5719):

@ShailKumar , @dan815 , @SyedaLovesPie

OpenStudy (anonymous):

9 moles of CO2 x (44 grams CO2 / 1 mole of CO2) = 396g of CO2

OpenStudy (anonymous):

So how are u going to fan me?

OpenStudy (anonymous):

The mass of 1 mole He is 4 gram. Calculate mass of 7 moles. Mass of 1 mole CO2 = 12+2*16 = 44 gram Now calculate mass of 9 moles

OpenStudy (hari5719):

lol ok ok

OpenStudy (anonymous):

Cheater

OpenStudy (anonymous):

You unfanned me then fanned me again